OpenStudy (anonymous):

Name the binomial you can multiply by (x + 9) to get the product x^2+5x-36 . A. 4 - x B. x + 6 C. x - 4 D. x - 6

OpenStudy (saifoo.khan):

Divide them. \[\frac{x^2 + 5x - 36}{x+9}\]Solve.

OpenStudy (anonymous):

thx :)

OpenStudy (saifoo.khan):

Np

OpenStudy (anonymous):

x2 - 5x - 36 = 0 (x - 9) (x + 4) = 0 x = 9 and x = -4 does this seem right?

OpenStudy (saifoo.khan):

Perfect.

OpenStudy (anonymous):

Thank you so much for helping me ...

OpenStudy (saifoo.khan):

No problem.